log(x^2-x-5)= 0
I canot figure out the way to solve this. I looked at the example in my book and it was not clear. Solve for x
my guess is to complete the square, but i am not sure how to do this

Answer by nerdybill(5404)

log(x^2-x-5)= 0
x^2-x-5= 10^0
x^2-x-5= 1
x^2-x-6= 0
(x-3)(x+2)=0
.
x = {-2, 3}